Find three consecutive integers such that twice the second subtracted from the sum of the other three integers is sixteen.

Answer by nerdybill(7384)

Find three consecutive integers such that twice the second subtracted from the sum of the other three integers is sixteen.
.
Let x = first integer
then
x+1 = second
x+2 = third
.
(x + x+1 + x+2)-2(x+1) = 16
(3x+3)-(2x+2) = 16
3x+3-2x-2 = 16
x+3-2 = 16
x+1 = 16
x = 15
.
Solution: 15, 16, and 17